# GPUDrive: Data-driven, multi-agent driving simulation at 1 million FPS

### v2

@@ Abstract
Multi-agent learning algorithms have been successful at generating superhuman planning in various games but have had limited impact on the design of deployed multi-agent planners. A key bottleneck in applying these techniques to multi-agent planning is that they require billions of steps of experience. To enable the study of multi-agent planning at scale, we present GPUDrive, a GPU-accelerated, multi-agent simulator built on top of the Madrona Game Engine that can generate over a million simulation steps per second. Observation, reward, and dynamics functions are written directly in C++, allowing users to define complex, heterogeneous agent behaviors that are lowered to high-performance CUDA. We show that using GPUDrive we can effectively train reinforcement learning agents over many scenes in the Waymo Open Motion Dataset, yielding highly effective goal-reaching agents in minutes for individual scenes and enabling agents to navigate thousands of scenarios within hours. The code base with pre-trained agents is available at https://github.com/Emerge-Lab/gpudrive.
@@ 1 Introduction
To address these challenges and unlock multi-agent learning as a tool for generating capable self-driving planners, we introduce GPUDrive. GPUDrive is a simulator intended to mix real-world driving data with simulation speeds that enable the application of sample-inefficient but effective RL algorithms to planner design. GPUDrive runs at over a million steps per second on both consumer-grade and datacenter-class GPUs and has a sufficiently light memory footprint to support hundreds to thousands of simultaneous worlds (environments) with hundreds of agents per world. GPUDrive supports the simulation of a variety of sensor modalities, from LIDAR to a human-like view cone, enabling GPUDrive to be used for studying the effects of different sensor types on resultant agent characteristics. Finally, GPUDrive takes in driving logs and maps from existing self-driving datasets, enabling the mixing of tools from imitation learning with reinforcement learning algorithms. This enables the study of both the development of autonomous vehicles and the learning of models of human driving, cycling, and walking behavior.
@@
  • We provide gymnasium (Towers et al., 2024) environments in both torch and jax that can be easily configured with standard open-source multi-agent RL and imitation learning libraries. Additionally, we have open-sourced a basic policy-gradient training loop that can be readily used to develop agents.
@@ 2 Related work
There are various open-source frameworks available that support hardware-accelerated reinforcement learning environments. These environments are generally written directly in an acceleration framework such as Numpy (Harris et al., 2020), Jax (Bradbury et al., 2018), or Pytorch (Ansel et al., 2024). In terms of multi-agent accelerated environments, standard benchmarks include JaxMARL(Rutherford et al., 2023), Jumanji (Bonnet et al., 2024), and VMAS (Bettini et al., 2022) which primarily feature fully cooperative or fully competitive tasks. Each benchmark requires the design of custom accelerated structures per environment. In contrast, GPUDrive focuses on a mixed motive setting and is built atop Madrona, an extensible ECS-based framework in C++, enabling GPU acceleration and parallelization across environments (Shacklett et al., 2023). Madrona comes with vectorization of key components of embodied simulation such as collision checking and sensors such as LIDAR. GPUDrive can support hundreds of controllable agents in more than 100,000 distinct scenarios, offering a distinct generalization challenge and scale relative to existing benchmarks. Moreover, GPUDrive includes a large dataset of human demonstrations, enabling imitation learning, inverse RL, and combined IL-RL approaches.
@@ 3.2 Simulator features
Figure 1: Four example scenarios from the Waymo Open Motion Dataset rendered in GPUDrive. The blue circles indicate vehicle destinations.
@@
GPUDrive supports a variety of observation spaces intended to enable heterogeneous types of agents. Fig. 2 depicts the three types of supported state spaces. The first mode is somewhat unphysical in which all agents and road objects within a fixed radius are observable to the agent. This mode is intended primarily for debugging and quick testing, enabling a user to minimize the amount of partial observability in the environment. The other two modes are based on a GPU-accelerated LIDAR scan, representing what an autonomous vehicle would be able to see and what a human would likely be able to see respectively. Both modes are based on casting LIDAR rays; to model human vision we simply restrict the LIDAR rays to emanate in a smaller, controllable-sized cone that can be rotated through an action corresponding to head rotation. Note that since all objects are represented as bounding boxes of fixed height, the LIDAR observations are over-conservative as in reality it is frequently possible to see over the hoods of cars as their height is lower than the body of the rest of the car.
@@
Note that this model does not factor in the length of the car, causing both long and short objects to have identical dynamics. However, computing the expert actions and then using them to mimic the expert trajectory under this model leads to lower tracking error than the default bicycle model.
@@
We terminate an agent’s episode when they achieve their goal position or collide. As it is not clear where an agent should go next after it reaches its goal, we simply remove it from the scene afterwards (we note that an alternative might be to generate a new goal for the agent to drive to). Car and cyclist agents are considered to be in collision when they drive through a road edge while pedestrian agents are allowed to cross road edges.
@@
  • Absence of a map. The current version of the simulator does not have a well-defined notion of lanes or a higher-level road map which makes it challenging for algorithmic approaches that require maps. The absence of this feature also makes it challenging to define rewards such as ”stay lane-centered.”
@@
  • Unsolvable goals. Due to incorrect labels of some road points in the Waymo dataset, such as an exit to a parking lot being labelled as an impassable road edge, some agent goals (roughly 2%) are unreachable. For these agents, we default them to simply replaying their logged trajectory and do not treat them as agents.
  • Variance in controllable agents per scenario. In most scenes, a significant fraction (25-75%) of the agents are not controllable because they do not move (e.g. parked cars). This is because the dataset is collected from an autonomous vehicle’s sensors, so some agents’ initial states are recorded after the first time-step of the simulator. These agents are not included to avoid collisions with other agents.
@@ 4 Simulator performance
The following Sections describe the simulator speed. Section 4.1 first shows the raw simulator speed and peak goodput (throughput achieved by the number valid number of agents in a scene). Section 4.2 then investigates the impact on reinforcement learning workflows by evaluating the time it takes to train reinforcement learning agents through Independent PPO (IPPO) (Yu et al., 2022), a widely used multi-agent learning algorithm. IPPO is implemented atop Stable Baselines 3 (Hill et al., 2018).
@@ 4.1 Simulation speed
Figure 3: Peak goodput of GPUDrive on a consumer-grade and datacenter-class GPU compared to original, CPU and GPU-based, implementations using the radial filter observation. Left: The total number of agent steps per second (ASPS) is the number of objects for which our system computes observations at each time step. To ensure a fair comparison, we align the conditions with those used in (Gulino et al., 2024), where all cars, bicyclists, and pedestrians are considered valid experience-generating agents. Center: The distribution of controllable agents across 512 scenarios in the Waymo Open Motion Dataset $\mu\approx 10.8\,(\text{red line}),\sigma\approx 9.3$). Right: The total number of controllable agent steps per second (CASPS) as we increase the number of worlds (parallelism).
@@ 4.2 End-to-end speed and performance
Figure 5 contrasts the number of steps (experience) and the corresponding time required to solve 10 scenarios from the WOMD between Nocturne (Vinitsky et al., 2022) and GPUDrive. We compare Nocturne to GPUDrive, as it is the closest simulator in functionality, except that it is implemented on the CPU. For benchmarking purposes, we say a scene is solved when agents can navigate to their designated target position $95\%$ of the time without colliding or going off-road. Ceteris paribus (Details in Appendix D), GPUDrive achieves a 30-40x training speedup, solving 10 scenarios in less than 15 minutes compared to approximately 10 hours in Nocturne.
Figure 5: From hours to minutes. Left: Training performance (goal-reaching rate) as a function of the global step (AFPS). Center: Training performance as a function of wall-clock time. Right: Comparison of the total time to solve the same 10 scenarios while replicating environmental and experimental conditions as closely as possible. Runs are averaged across three seeds, see Appendix D for the hyperparameters and training details. The green dotted line indicates optimal performance.
@@ 5 Conclusion
In this work, we present GPUDrive, a GPU-accelerated, data-driven simulator. GPUDrive is intended to help generate the billions of samples that are likely needed to achieve effective reinforcement learning for multi-agent driving planners. By building atop the Madrona Engine (Shacklett et al., 2023), we can scale GPUDrive to hundreds of worlds with potentially thousands of agents leading to throughput of millions of steps per second. This throughput occurs while synthesizing complex observations such as LiDAR. We show that this throughput has consequent implications for training reinforcement learning agents, leading to the ability to train agents to solve any particular scene in minutes and in seconds when amortized across many scenes. We release the simulator and integrated trained agents to enable further research.
This paper is a first step in scaling up reinforcement learning for multi-agent planning in safety-critical, mixed human-autonomous settings. However, several important challenges remain for future work. Firstly, we have not yet identified the optimal hyperparameter settings to effectively utilize the collected data, resulting in training being the bottleneck instead of data collection time. Secondly, while the simulator is fast, collecting data for reinforcement learning leads to frequent reset calls that significantly impact throughput. Lastly, training fully human-level drivers in the simulator to navigate without crashing in any scenario remains an unresolved challenge. We have shown that our simulator allows training of driving agents to achieve a 95% success rate in a few hours. However, this still does not match human capabilities or the safety standards we strive for. We hope GPUDrive will help the community tackle this open challenge.
@@ Appendix A Reproducibility
A.1 Code Reproducibility
All code required to reproduce the paper is open-sourced at
https://github.com/Emerge-Lab/gpudrive under release number v0.2.0.
A.2 Computational Resources
All RL experiments in this paper were run on an NVIDIA RTX 8000 or A100. Total resources for the paper correspond to less than 2 GPU days.
@@ Appendix C License Details and Accessibility
Our code is released under an MIT License. The Waymo Motion dataset is released under a Apache License 2.0. The code is available at <ANONYMIZED-DURING-REVIEW>.
@@ D.1 End-to-end performance
Goal achievement rate
Vehicle collisions
Off-road events
Figure 7: Key performance metrics as a function of training time, grouped by the number of unique scenes in a batch, as reported in Figure 6. Left: The aggregate percentage of agents that achieved their goal. Center: The aggregate percentage of agents that collided with another vehicle. Right: The aggregate number of vehicles that crossed a road edge.

### v3

@@ Abstract
Multi-agent learning algorithms have been successful at generating superhuman planning in various games but have had limited impact on the design of deployed multi-agent planners. A key bottleneck in applying these techniques to multi-agent planning is that they require billions of steps of experience. To enable the study of multi-agent planning at scale, we present GPUDrive. GPUDrive is a GPU-accelerated, multi-agent simulator built on top of the Madrona Game Engine capable of generating over a million simulation steps per second. Observation, reward, and dynamics functions are written directly in C++, allowing users to define complex, heterogeneous agent behaviors that are lowered to high-performance CUDA. Despite these low-level optimizations, GPUDrive is fully accessible through Python, offering a seamless and efficient workflow for multi-agent, closed-loop simulation. Using GPUDrive, we train reinforcement learning agents on the Waymo Open Motion Dataset, achieving efficient goal-reaching in minutes and scaling to thousands of scenarios in hours. We open-source the code and pre-trained agents at
www.github.com/Emerge-Lab/gpudrive
@@ 1 Introduction
Figure 1: Extremely fast multi-agent simulation with GPUDrive. Top: Bird’s-eye view of Waymo Open Motion Dataset scenarios in GPUDrive, with boxes marking controlled agents and circles denoting their goals. Bottom: Corresponding agent views, centered on one agent. Observations can be easily configured based on the user’s objectives. Here, agents are provided with a scene view through a relative coordinate frame. Shown are nearby road points within a configurable radius (set to 50 meters) and the relative positions of other agents in the scene.
@@
To address these challenges and unlock multi-agent learning as a tool for generating capable self-driving planners, we introduce GPUDrive. GPUDrive is a simulator intended to mix real-world driving data with simulation speeds that enable the application of sample-inefficient but effective RL algorithms to the design of autonomous planners. GPUDrive runs at over a million steps per second on both consumer-grade and datacenter-class GPUs and has a sufficiently light memory footprint to support hundreds to thousands of simultaneous worlds (environments) with hundreds of agents per world. GPUDrive supports the simulation of a variety of sensor modalities, from LIDAR to a human-like view cone, enabling GPUDrive to be used for studying the effects of different sensor types on resultant agent characteristics. Finally, GPUDrive takes in driving logs and maps from existing self-driving datasets, enabling the mixing of tools from imitation learning with reinforcement learning algorithms. This enables the study of both the development of autonomous vehicles and the learning of models of human driving, cycling, and walking behavior.
@@
  • We provide gymnasium (Towers et al., 2024) environments in both torch and jax that can be easily configured with standard open-source multi-agent RL and imitation learning libraries. Additionally, we open-sourced two policy-gradient training loops that can be readily used to develop agents.
@@ 2 Related work
There are various open-source frameworks available that support hardware-accelerated reinforcement learning environments. These environments are generally written directly in an acceleration framework such as Numpy (Harris et al., 2020), Jax (Bradbury et al., 2018), or Pytorch (Ansel et al., 2024; Makoviychuk et al., 2021; Panerati et al., 2021). In terms of multi-agent accelerated environments, standard benchmarks include JaxMARL(Rutherford et al., 2023), Jumanji (Bonnet et al., 2024), VMAS (Bettini et al., 2022) Gigastep  (Lechner et al., 2024) which primarily feature fully cooperative or fully competitive tasks. Each benchmark requires the design of custom accelerated structures per environment. In contrast, GPUDrive focuses on a mixed motive setting and is built atop Madrona, an extensible ECS-based framework in C++, enabling GPU acceleration and parallelization across environments (Shacklett et al., 2023). Madrona comes with vectorization of key components of embodied simulation such as collision checking and sensors such as LIDAR. GPUDrive can support hundreds of controllable agents in more than 100,000 distinct scenarios, offering a distinct generalization challenge and scale relative to existing benchmarks. Moreover, GPUDrive includes a large dataset of human demonstrations, enabling imitation learning, inverse RL, and combined IL-RL approaches.
@@ 3.2 Simulator features
Figure 2: Example scenarios from the Waymo Open Motion Dataset rendered in GPUDrive. The blue boxes and circles indicate agents and their respective destinations.
@@
GPUDrive supports a variety of observation spaces intended to enable heterogeneous types of agents. Fig. 7 depicts the three types of supported state spaces. The first mode is somewhat unphysical in which all agents and road objects within a fixed radius are observable to the agent. This mode is intended primarily for debugging and quick testing, enabling a user to minimize the amount of partial observability in the environment. The other two modes are based on a GPU-accelerated LIDAR scan, representing what an autonomous vehicle would be able to see and what a human would likely be able to see respectively. Both modes are based on casting LIDAR rays; to model human vision we simply restrict the LIDAR rays to emanate in a smaller, controllable-sized cone that can be rotated through an action corresponding to head rotation. Note that since all objects are represented as bounding boxes of fixed height, the LIDAR observations are over-conservative as, in reality, it is frequently possible to see over the hoods of cars as their height is lower than the body of the rest of the car.
@@
Note that this model does not factor in the length of the car, causing both long and short objects to have identical dynamics. However, we have observed that computing the expert actions and then using them to mimic the expert trajectory under this model leads to lower tracking error than the default bicycle model.
@@
We terminate an agent’s episode when they achieve their goal position and support an option to also terminate the episode if the agent collides. As it is not clear where an agent should go next after it reaches its goal, we simply remove it from the scene afterwards (we note that an alternative might be to generate a new goal for the agent to drive to). Car and cyclist agents are considered to be in collision when they drive through a road edge while pedestrian agents are allowed to cross road edges.
@@
  • Absence of a map. The current version of the simulator does not have a well-defined notion of lanes or a higher-level road map which makes it challenging for algorithmic approaches that require maps. The absence of this feature also makes it challenging to define rewards such as "stay lane-centered."
@@
  • Unsolvable goals. Due to incorrect labels of some road points in the Waymo dataset, such as an exit to a parking lot being labeled as an impassable road edge, some agent goals (roughly 2%) are unreachable. For these agents, we default them to simply replaying their logged trajectory and do not treat them as agents.
  • Initialization modes: In many scenarios, a significant portion of agents (25-75%) are already at or very close to their goal positions, such as parked cars. To highlight the difficulty of controlling agents, GPUDrive supports different initialization modes. By default, the initialization mode is set to “all nontrivial” meaning that only agents more than 2 meters away from their target positions are considered controllable. Note that while this sharply reduces the number of agents in the scene, it more accurately represents actual driving challenges.
@@ 4 Simulator performance
The following Sections describe the simulator speed. Section 4.1 first shows the raw simulator speed and peak goodput (throughput achieved by the valid number of agents in a scene). Section 4.2 then investigates the impact on reinforcement learning workflows by evaluating the time it takes to train reinforcement learning agents through Independent PPO (IPPO) (Yu et al., 2022), a widely used multi-agent learning algorithm. The performance results here are from IPPO implemented with PufferLib (Suarez, 2024). A slower IPPO implementation based on Stable Baselines 3 (Hill et al., 2018) is also available in the repo.
@@ 4.1 Simulation speed
Figure 3: Peak goodput of GPUDrive on a consumer-grade and datacenter-class GPU compared to original, CPU and GPU-based, implementations using the radial filter observation. Left: The total number of agent steps per second (ASPS) is the number of objects for which our system computes observations at each time step. To ensure a fair comparison, we align the conditions with those used in (Gulino et al., 2024), where all cars, bicyclists, and pedestrians are considered valid experience-generating agents. Center: The distribution of controllable agents across 512 scenarios in the Waymo Open Motion Dataset $\mu\approx 10.8\,(\text{red line}),\sigma\approx 9.3$). These numbers are obtained using the ”nontrivial“ initialization mode in GPUDrive, which initializes only the agents that are more than 2 meters away from their final position. Right: The total number of controllable agent steps per second (CASPS) as we increase the number of worlds (parallelism).
@@ 4.2 End-to-end speed and performance
Figure 5 contrasts the number of steps (experience) and the corresponding time required to solve 10 scenarios from the WOMD between Nocturne (Vinitsky et al., 2022) and GPUDrive. We compare Nocturne to GPUDrive, as it is the closest simulator in functionality, except that it is implemented on the CPU. For benchmarking purposes, we mark a scene as solved when agents can navigate to their designated target position $95\%$ of the time without colliding or going off-road. Ceteris paribus (Details in Appendix D), GPUDrive achieves a 200 - 300x training speedup, solving 10 scenarios in less than 3 minutes compared to approximately 10 hours in Nocturne.
In absolute terms, the number of controlled agent steps per second for the Pufferlib PPO implementation ranges from 200K to 500K, depending on the hardware used and factors such as the number of visible road points per agent.
Figure 5: From hours to seconds. Left: Training performance (goal-reaching rate) as a function of the global step of the controlled agents (CASPS). Center: Training performance as a function of wall-clock time in seconds. Right: Training performance as a function of wall-clock time where the x-axis is on a log scale. Runs are averaged across three seeds, replicating environmental and experimental conditions as closely as possible. See Appendix D for the hyperparameters and training details. The green dotted line marks optimal performance (all agents reach their goal without collisions).
@@ 5 Conclusion
In this work, we present GPUDrive, a GPU-accelerated, multi-agent, and data-driven simulator. GPUDrive is intended to help generate the billions of samples that are likely needed to achieve effective reinforcement learning for multi-agent driving planners. By building atop the Madrona Engine (Shacklett et al., 2023), we can scale GPUDrive to hundreds of worlds with potentially thousands of agents leading to throughput of millions of steps per second. This throughput occurs while synthesizing complex observations such as LiDAR. We show that this throughput has consequent implications for training reinforcement learning agents, leading to the ability to train agents to solve any particular scene in minutes and in seconds when amortized across many scenes. We release the simulator and integrated trained agents to enable further research.
Future work and simulator extensions.
This paper represents an initial step toward scaling reinforcement learning for multi-agent planning in safety-critical, mixed human-autonomous settings. Several important opportunities remain for future work:
  • Agent performance: Training agents to navigate without crashing in any scenario, matching human capabilities, remains an unresolved challenge. Currently, agents trained in GPUDrive achieve a 95% success rate in a few hours, but this still falls short of the standards we aim for. Getting to 100% goal reaching performance is left for future work.
  • Integrating multiple datasets: Future work will focus on integrating various datasets, such as NuPlan or NuScenes (Caesar et al., 2020). This integration will enable training on data from multiple geographic locations, including Boston and Singapore. To achieve this, we need to establish a data-processing pipeline that ensures the datasets are converted into a format compatible with the simulator.
  • Diverse realistic sim agents: We plan to extend GPUDrive by introducing a variety of simulation agents that represent a broad range of human-like behaviors. This will improve sim realism.
6 Reproducibility Statement
The simulator’s source code and data are available at https://github.com/Emerge-Lab/gpudrive. To ease reproducibility, we provide Dockerfiles to simplify setup. The experiments in the paper can be reproduced using a single file on a single A100 in 16 hours or slightly longer for less performant hardware. The data has been pulled from the Waymo Motion dataset and parsed into JSON files that are used to initialize the simulator; these are all available from Huggingface datasets.
7 Ethics Statement
This paper presents GPUDrive, a GPU-accelerated simulator for multi-agent learning in autonomous driving. We use publicly available datasets, such as the Waymo Open Motion Dataset (Ettinger et al., 2021), which are anonymized to protect privacy. GPUDrive is intended for research purposes and not for real-world deployment without further validation. We recognize the risks of autonomous systems and emphasize the importance of safety and fairness in their application. Although this work does not involve human participants directly, it leverages real-world data that may reflect human behavior, and we strive to avoid harm or discrimination. The codebase, pre-trained agents, and evaluation scripts are openly available to promote transparency. There are no conflicts of interest or external sponsorships affecting this research. We are committed to ethical practices in data handling, model deployment, and research integrity.
Acknowledgments
This work is funded by the C2SMARTER Center through a grant from the U.S. DOT’s University Transportation Center Program. The contents of this report reflect the views of the authors, who are responsible for the facts and the accuracy of the information presented herein. The U.S. Government assumes no liability for the contents or use thereof. This work was also supported in part through the NYU IT High-Performance Computing resources, services, and staff expertise.
@@ Appendix A Reproducibility
A.1 Extra figures
A.2 Code Reproducibility
All code required to reproduce the paper is open-sourced at https://github.com/Emerge-Lab/gpudrive/tree/main under release number v0.4.0
A.3 Computational Resources
All RL experiments in this paper were run on an NVIDIA RTX 4080 or A100. Total resources for the paper correspond to less than 2 GPU days.
@@ Appendix C License Details and Accessibility
Our code is released under an MIT License. The Waymo Motion dataset is released under a Apache License 2.0. The code is available at https://github.com/Emerge-Lab/gpudrive.
@@ D.1 End-to-end performance
Figure 8: Key performance metrics as a function of training time grouped by the number of unique scenes in a batch reported in Figure 6. Left: The aggregate percentage of agents that achieved their goal. Center: The aggregate percentage of agents that collided with another vehicle. Right: The aggregate number of vehicles that crossed a road edge.